# Art or Artifice? Large Language Models and the False Promise of Creativity

### v1

@@ 1. Introduction
Recent work has explored the potential of using large language models (LLMs) as assistants in creative writing tasks, from stories to screenplays (Mirowski et al., 2023; Yang et al., 2022; Lee et al., 2022; Ippolito et al., 2022; Chen et al., 2023b). An important aspect of this research is to investigate LLMs’ capabilities in terms of their ability to both generate creative content and assess whether a piece of writing is creative, with the ultimate goal of informing interaction design. However, evaluating objectively the creativity of a piece of writing is challenging.
@@
We propose a protocol to evaluate creativity as product grounded in a widely accepted protocol that evaluates creativity as process — the Torrance Tests of Creative Thinking (TTCT) (Torrance, 1966). Based on Guilford’s work on divergent thinking (Guilford, 1967), TTCT measures creativity as a process by testing participants’ abilities in dealing with unusual uses of objects, specific situations, or impossibilities. TTCT is centered around evaluating four dimensions of creativity: fluency (the sheer volume of meaningful ideas produced in reaction to a given stimulus), flexibility (the diversity of categories within the responses), originality (the uniqueness or novelty of answers) and elaboration (the depth or granularity of details within the responses). While the direct application of TTCT might not be possible across diverse creative domains (Amabile, 1982; Baer and McKool, 2009), its four fundamental dimensions have proven adaptable (Trisnayanti et al., 2019; McIntyre et al., 2003; Bourdeau et al., 2020). Thus, based on TTCT and using the Consensual Assessment Technique (CAT) (Amabile, 1982) we design the Torrance Tests for Creative Writing (TTCW) to evaluate creativity as product (Section 3 Design Principle 1 and 2; Figure 1). The Consensual Assessment Technique states that the most valid assessment of the creativity of an idea or creation in any field is the collective judgment of experts in that field. Thus, to design TTCW, we asked 8 creative writing experts in a formative study (Section 4) to propose creativity measures for the evaluation of short fictional stories aligned with the four Torrance dimensions. This resulted in 14 binary tests organized across the four original Torrance dimensions of fluency, flexibility, originality and elaboration (Section 4.2).
Figure 1. Pipeline showing the construction of TTCW and evaluation of short stories using the TTCW framework
@@
To empirically validate TTCW as an evaluation protocol for creativity as product (fictional short stories), we build a benchmark consisting of 48 short stories: 12 stories written by professionals, and 36 by three top performing LLMs (ChatGPT (OpenAI, 2022), GPT4 (OpenAI, 2023) and Claude 1.3 (Anthropic, 2022)) with 1400 words on average per story (Section 5; Figure 1). We recruit a new set of 10 experts, again adhering to CAT, to administer the 14 tests of TTCW on each story, collecting 3 evaluations per story. In this TTCW implementation with experts as assessors we aim to answer three research questions:
@@ 3. RQ3:
Prior work has argued that even though LLMs might not be able to directly generate a creative piece of writing, they might be utilized in providing feedback to authors during their writing process (Ippolito et al., 2022). Thus, we perform a study of TTCW implementation with LLMs as assessors (Section  6; and Figure 1) to understand whether LLMs can be used to assess creative writing. We expanded each test into a detailed prompt, and measured whether LLM assessments correlate with collected expert judgments. Our analysis reveals that for the most part, LLMs are not capable of administering the TTCW tests, as the three LLMs we experiment with achieve correlations with experts that are close to zero.
@@
  • We study the abilities of LLMs to generate stories that pass/fail the TTCW tests and their ability to reliably assess the creativity of stories following the TTCW framework through correlation with human judgments. Our findings show that LLM-generated stories are three to ten times less likely to pass TTCW tests compared to expert-written stories, as well as the fact that current state-of-the-art LLMs are not yet capable of reproducing expert assessments when administering TTCW tests. To enable future research in this fast-evolving domain, we release the large-scale annotation of 2,000+ TTCW assessments [^1] https://github.com/salesforce/creativity_eval, each accompanied with a natural language expert explanation.
@@ 5. TTCW Implementation with Experts as Assessors
In this Section, we detail our implementation of creative writing evaluation using the TTCW framework derived previously. We first carefully select the 48-short stories included in the evaluation, then go over the 2.5-hour study design protocol that our 10 experiment participants followed, and finally analyze the results based on the 2,000+ individual tests administered by the experts.We leverage the collected annotations to answer the following research questions:
@@ 5.1. Data Selection
Prior studies in the creativity of model-generated text have shown that technologies such as LLMs are capable of generative long and coherent stories (Yang et al., 2023, 2022), as well as act as assist and collaborate with creative writers (Yuan et al., 2022; Ippolito et al., 2022; Mirowski et al., 2023). For practical considerations, these studies limit their evaluation solely to model-generated content, typically only claiming relative improvements from one system to another, and do not establish whether a gap remains between high-quality human-written stories and model-generated text. Our research employs a rigorous evaluation protocol that juxtaposes human-written short stories against those constructed by LLMs to discern the potential gap, if any, in their creative quality. The dataset selection process is visually summarized in Figure 2. We first collect 12 short stories from The New Yorker collection of short stories  [^10] https://www.newyorker.com/tag/short-stories. Our selection criterion involved choosing short stories with diverse authors and plots. Spanning from August 13, 2020, to May 8, 2023, these stories (ranging between 1000 to 2,400 words as delineated in Figure 5 in the Appendix) include compositions from acclaimed authors such as Haruki Murakami to Nobel laureate Annie Ernaux. The titles and a one-sentence plot summary (generated by GPT-4 and verified by humans) of included New Yorker short stories are listed in Table  10 in Appendix.
@@ 5.3. Participant Recruitment
To test the robustness and validity of TTCW-based evaluation, we chose to recruit a new set of experts to conduct the evaluation and not re-hire the ones from our formative study that played a role in the creation of the tests. We posit that such a choice demonstrates the fact that the TTCW can be administered by any knowledgeable expert provided solely with the tests and their expanded explanations. We recruited 10 participants on the User Interviews platform and listed their background in Table 4. Four of these participants are associated with the creative writing departments at leading American academic institutions, with considerable experience conducting undergraduate and graduate-level courses. Two participants function as literary agents at a top-tier, full-service US literary agency representing well-recognized authors, and four are professional writers with a Master of Fine Arts in Fiction or Poetry.
To experimentally analyze the reproducibility and validity of the TTCW, we randomly assigned each story group to 3 distinct experts. This allows us to study the agreement levels between experts on individual tests as well as in aggregate. After each task, an expert participant was given the option to receive another group of stories, and our participants completed on average 3.6 tasks over a period of 3 weeks, for a total of 36 assessments (i.e., 3 for each of the twelve groups). Because each assessment contains four stories, and each story was evaluated using the 14 TTCW, we collected a total of 2,016 binary labels and expert-written justifications for these labels.
5.4. Results
@@ 5.4.2. Reproducibility of TTCW
Since the TTCW are designed to be additive, we further compute an aggregate score for each story by counting the number of TTCW tests a story passes. We visualize the results of this aggregate measure in Figure  3. Since the aggregate measure is numerical (ranges from 0 to 14), we use Pearson correlation to measure agreement among experts. At this aggregate level, we obtain a correlation($\rho$) of 0.69, showing strong agreement among experts on the number of tests a story passes. In other words, even though experts reach slightly lower agreement on which exact TTCW a story passes or fails, they achieve strong agreement on the number of tests a story passes overall. This experimental finding confirms the importance of Design Principle 4, and the need for the tests to be performed as a set to achieve a reproducible evaluation of creativity for a given short story. When the objective is to evaluate the broad creativity in a short story, it is our recommendation that all fourteen tests should be administered as a set by one expert annotator, rather than by different experts or administering only an individual test, as this increases reproducibility of results.
@@ 5.4.3. Comparative Evaluation Results
Looking at the ranking results, human-written New Yorker stories were ranked as the most preferred story 89% of the time, while the GPT-3.5-generated stories rank as least preferred roughly two-thirds of the time. When comparing between GPT-4 and Claude, Claude is almost twice as likely to rank as second (behind the human-written story) and was the most preferred on three of the four assessments in which the New Yorker story was not chosen as the most favored. These ranking results confirm and accentuate the observation from the test passing rates analysis that Claude V1.3 generates higher-quality short stories than models in the GPT family.
@@
The attribution results paint a similar picture, with New Yorker stories predominantly attributed to an experienced writer, while LLM-generated stories get attributed to AI or an amateur writer. Interestingly, Claude V1.3 is more likely to be attributed to an amateur writer than an AI, whereas GPT3.5 and GPT4 stories are 80%+ attributed to AI. One hypothesis for such a behavior could be that the participants in our study might are more familiar with text written by OpenAI models, as these models are commercially more successful, providing an element of surprise to Claude-written text.
6. TTCW Implementation with LLMs as Assessors
Expert annotation such as the one we perform in Section 5 is costly: based on our evaluation protocol, evaluating a 1500-2500 word short story with a qualified expert costs \$20, and requires roughly 30 minutes of the expert’s time. Prior work has shown the promise of using LLMs in text evaluation. For instance, GPT3.5 and GPT4 are effective at evaluating the factual consistency of a summary to its document (Laban et al., 2023), or measuring the coherence of a summary (Gao et al., 2023). GPTEval (Liu et al., 2023) employs the framework of using large language models with chain-of-thoughts (CoT) (Wei et al., 2022) to assess the quality of NLG outputs. Recent work has also applied LLM-based evaluation to the creative domain (Rajani et al., 2023), claiming that GPT4 can achieve a high correlation with humans when evaluating brainstorming or creative generation tasks. In this section we describe our TTCW implementation with LLMs as assessor to understand LLMs ability to assess creative writing.
Prior work Wei et al. (2022) has shown how generating a chain of thought – a series of intermediate reasoning steps – significantly improves the ability of large language models to perform complex reasoning. Taking advantage of this we design the prompts/instructions for large language models in a slightly different fashion than for the human experts as can be seen in Table 3(Row 3 (Human Instruction) vs Row 4 LLM Instruction). To help the model make an informed decision we first ask it to list out elements specific to any given test such as “elements in the story that call to each of the five senses” for the World Building and setting test followed by asking it to decide overall and then provide its reasoning before choosing an answer between ‘Yes’ or ‘No’. The exact prompt contains (1) the story, (2) the expanded TTCW context, (3) the TTCW question, and (4) a LLM-specific instruction guiding the model to perform the task in a chain-of-though manner. We then measure model agreement with the majority vote of the three experts that conducted the test, using Cohen’s Kappa.
6.1. Results
Dimension Test GPT3.5 GPT4 Claude
Fluency Understandability & Coherence -0.01 -0.01 -0.17
Narrative Pacing 0.05 0.0 -0.22
Scene vs Exposition -0.03 -0.08 -0.23
Literary Devices & Language Proficiency
0.04 -0.09 -0.11
Narrative Ending -0.02 0.02 0.02
Flexibility Emotional Flexibility -0.04 0.0 0.09
Perspective & Voice Flexibility 0.0 0.26 0.14
Structural Flexibility -0.04 0.0 -0.07
Originality Originality in Form 0.08 0.09 0.03
Originality in Thought 0.19 0.31 0.15
Originality in Theme & Content 0.06 -0.01 0.18
Elaboration World Building & Setting 0.0 0.00 0.09
Character Development -0.08 0.02 0.00
Rhetorical Complexity 0.0 0.0 0.02
Average 0.016 0.035 -0.006
7. Discussion
In an optional exchange with the expert participants who participated in the annotation of Section 5, participants were given the opportunity to describe how they differentiated between AI-generated and human-written stories. Table 14 in the Appendix lists the replies of five experts, which we color-coded to highlight the recurrent issues that led them to believe that a story is written by an AI. In particular, E5, E4, E1 thought AI struggles at Narrative Ending. E5 and E4 highlighted that AI-written stories would forestall the ending by getting bigger in scope. E1 highlighted that AI-written stories would have multiple disparate endings. E5, E4, E1, and E2 all highlighted that AI-written stories would often contain abstruse and incoherent metaphors that do not add meaning or extremely cliched or simple metaphors thereby demonstrating poor Language Proficiency and use of Literary Devices. E1 highlighted one such example in a story - However, she managed to laugh louder and louder until her laughter transformed into an embrace of the sun’s atmosphere.
E1 and E2 further highlighted that characters in AI-written stories have poor Rhetorical Complexity and are often lacking in subtext. E1 further added that AI-written stories operate in a nearly opposite and Drax-like fashion in which there is only literal meaning, and that literal meaning is often nonsensical, or at least presented without any of the context that might make it seem like something a human would say or do. E1 highlighted an exchange below from an AI-written story

### v2

@@ 1. Introduction
Recent work has explored the potential of using large language models (LLMs) as assistants in creative writing tasks, from stories to screenplays (Mirowski et al., 2023; Yang et al., 2022; Lee et al., 2022; Ippolito et al., 2022; Chen et al., 2023c). An important aspect of this research is to investigate LLMs’ capabilities in terms of their ability to both generate creative content and assess whether a piece of writing is creative, with the ultimate goal of informing interaction design. However, evaluating objectively the creativity of a piece of writing is challenging.
@@
We propose a protocol to evaluate creativity as product grounded in a widely accepted protocol that evaluates creativity as process — the Torrance Tests of Creative Thinking (TTCT) (Torrance, 1966). Based on Guilford’s work on divergent thinking (Guilford, 1967), TTCT measures creativity as a process by testing participants’ abilities in dealing with unusual uses of objects, specific situations, or impossibilities. TTCT is centered around evaluating four dimensions of creativity: fluency (the sheer volume of meaningful ideas produced in reaction to a given stimulus), flexibility (the diversity of categories within the responses), originality (the uniqueness or novelty of answers) and elaboration (the depth or granularity of details within the responses). While the direct application of TTCT might not be possible across diverse creative domains (Amabile, 1982; Baer and McKool, 2009), its four fundamental dimensions have proven adaptable (Trisnayanti et al., 2019; McIntyre et al., 2003; Bourdeau et al., 2020). Thus, based on TTCT and using the Consensual Assessment Technique (CAT) (Amabile, 1982) we design the Torrance Tests for Creative Writing (TTCW) to evaluate creativity as product (Section 3 Design Principle 1 and 2; Figure 1 Step 1). The Consensual Assessment Technique states that the most valid assessment of the creativity of an idea or creation in any field is the collective judgment of experts in that field. Thus, to design TTCW, we asked 8 creative writing experts in a formative study (Section 4) to propose creativity measures for the evaluation of short fictional stories aligned with the four Torrance dimensions. This resulted in 14 binary tests organized across the four original Torrance dimensions of fluency, flexibility, originality and elaboration (Section 4.2).
Figure 1. Pipeline showing the construction of TTCW and evaluation of short stories using the TTCW framework where Step 1) shows how experts leverage the process-oriented Torrance Test of Creative Thinking to create 14 tests for evaluating creativity in short stories as a product. Step 2) demonstrates artifact-centric testing where 4 stories based on a single plot are used as a product of creativity evaluation Step3) shows an evaluation of stories using the TTCW framework by both expert humans and LLMs where they each provide Yes/No answers to individual tests followed by natural language rationales justifying their decision.
@@
To empirically validate TTCW as an evaluation protocol for creativity as product (fictional short stories), we build a benchmark consisting of 48 short stories: 12 stories written by professionals, and 36 by three top performing LLMs (ChatGPT (OpenAI, 2022), GPT4 (OpenAI, 2023) and Claude 1.3 (Anthropic, 2022)) with 1400 words on average per story (Section 5; Figure 1 Step 2). We recruit a new set of 10 experts, again adhering to CAT, to administer the 14 tests of TTCW on each story, collecting 3 evaluations per story. In this TTCW implementation with experts as assessors (Figure 1 Step 3) we aim to answer three research questions:
@@ 3. RQ3:
Prior work (Ippolito et al., 2022) has argued that given the nature of LLMs where they regurgitate text seen during pre-training, they are often unable to directly generate a truly original and creative piece of writing (Also see Section 5.6). However, they might be utilized in providing feedback to authors during their writing process (Chakrabarty et al., 2023). Thus, we perform a study of TTCW implementation with LLMs as assessors (Section  6; and Figure 1 Step 3) to understand whether LLMs can be used to assess creative writing. We expanded each test into a detailed prompt, and measured whether LLM assessments correlate with collected expert judgments. Our analysis reveals that for the most part, LLMs are not capable of administering the TTCW tests, as the three LLMs we experiment with achieve correlations with experts that are close to zero.
@@
  • We study the abilities of LLMs to generate stories that pass/fail the TTCW tests and their ability to reliably assess the creativity of stories following the TTCW framework through correlation with human judgments. Our findings show that LLM-generated stories are three to ten times less likely to pass TTCW tests compared to expert-written stories, as well as the fact that current state-of-the-art LLMs are not yet capable of reproducing expert assessments when administering TTCW tests. To enable future research in this fast-evolving domain, we release the large-scale annotation of 2,000+ TTCW assessments, each accompanied with a natural language expert explanation.
@@ 5. TTCW Implementation with Experts as Assessors
In this Section, we detail our implementation of creative writing evaluation using the TTCW framework derived previously. We first carefully select the 48 short stories included in the evaluation, then go over the 2.5-hour study design protocol that our 10 experiment participants followed, and finally analyze the results based on the 2,000+ individual tests administered by the experts. We leverage the collected annotations to answer the following research questions:
@@ 5.1. Data Selection
Prior studies in the creativity of model-generated text have shown that technologies such as LLMs are capable of generative long and coherent stories (Yang et al., 2023, 2022), as well as act as assist and collaborate with creative writers (Yuan et al., 2022; Ippolito et al., 2022; Mirowski et al., 2023). For practical considerations, these studies limit their evaluation solely to model-generated content, typically only claiming relative improvements from one system to another, and do not establish whether a gap remains between high-quality human-written stories and model-generated text. Our research employs a rigorous evaluation protocol that juxtaposes human-written short stories against those constructed by LLMs to discern the potential gap, if any, in their creative quality. The dataset selection process is visually summarized in Figure 2. We first collect 12 short stories from The New Yorker collection of short stories  [^9] https://www.newyorker.com/tag/short-stories. Our selection criterion involved choosing short stories with diverse authors and plots. Spanning from August 13, 2020, to May 8, 2023, these stories (ranging between 1000 to 2,400 words as delineated in Figure 5 in the Appendix) include compositions from acclaimed authors such as Haruki Murakami to Nobel laureate Annie Ernaux. The titles and a one-sentence plot summary (generated by GPT-4 and verified by humans) of included New Yorker short stories are listed in Table  11 in Appendix.
@@
In our experiments, the stories either originate from experts or LLMs, and we do not include stories written by non-experts or ‘amateurs’. Although this was a consideration in the initial phases of the project, collecting amateur-written stories turned out to be a challenging and expensive process. Given the way our experiments are designed, it would require recruiting 12 non-experts to write stories given a plot since each expert-written story comes from a distinct author. The definition of an ‘expert’ or ‘amateur’ creative writer is difficult in a field that has unclear professional delineations. Collecting expert-written stories is straightforward as we rely on publication venue as a signal but identifying amateur writers is in itself a challenging task and might entail certain selection biases. A story written by any individual otherwise deemed as ‘amateur’ can still be of considerable high quality. Another alternative we considered is resorting to crowd-working platforms to recruit amateurs for writing stories but a recent study by Veselovsky et al. (2023) highlighted the fact that approximately 33-46% of crowd workers on such platforms currently utilize large language models (LLMs) to complete any assigned task, which would negatively impact our experiments’ reproducibility. Finally adding amateur-written stories would require increased budgets for evaluation. These difficulties govern our decision not to include amateur-written stories.
@@ 5.2.1. Expanded Expert Measure and Prompt Design
5.2.2. Expert Evaluation Protocol
  1. (1)
  2. (2)
  3. (3)
  4. (4)
5.3. Participant Recruitment
To test the robustness and validity of TTCW-based evaluation, we chose to recruit a new set of experts to conduct the evaluation and not re-hire the ones from our formative study that played a role in the creation of the tests. We posit that such a choice demonstrates the fact that the TTCW can be administered by any expert provided solely with the tests and their expanded explanations. We recruited 10 participants on the User Interviews platform and listed their background in Table 4. Four of these participants are associated with the creative writing departments at leading American academic institutions, with considerable experience conducting undergraduate and graduate-level courses. Two participants function as literary agents at a top-tier, full-service US literary agency representing well-recognized authors, and four are professional writers with a Master of Fine Arts in Fiction or Poetry.
To experimentally analyze the reproducibility and validity of the TTCW, we randomly assigned each story group to 3 distinct experts. This allows us to study the agreement levels between experts on individual tests as well as in aggregate. After each task, an expert participant was given the option to receive another group of stories, and our participants completed on average 3.6 tasks over 3 weeks, for a total of 36 assessments (i.e., 3 for each of the twelve groups). Because each assessment contains four stories, and each story was evaluated using the 14 TTCW, we collected a total of 2,016 binary labels and expert-written justifications for these labels.
5.4. Results
@@ 5.4.2. Reproducibility of TTCW
Since the TTCW are designed to be additive, we further compute an aggregate score for each story by counting the number of TTCW tests a story passes. We visualize the results of this aggregate measure in Figure  3. Since the aggregate measure is numerical (ranges from 0 to 14), we use Pearson correlation to measure agreement among experts. At this aggregate level, we obtain a correlation($\rho$) of 0.69, showing strong agreement among experts on the number of tests a story passes. In other words, even though experts reach slightly lower agreement on which exact TTCW a story passes or fails, they achieve strong agreement on the number of tests a story passes overall. This experimental finding confirms the importance of Design Principle 4, and the need for the tests to be performed as a set to achieve a reproducible evaluation of creativity for a given short story. When the objective is to evaluate the broad creativity in a short story, we recommend that all fourteen tests be administered as a set by one expert annotator, rather than by different experts or administering only an individual test, as this increases the reproducibility of the results.
@@ 5.4.3. Comparative Evaluation Results
Looking at the ranking results, human-written New Yorker stories were ranked as the most preferred story 89% of the time, while the GPT-3.5-generated stories ranked as least preferred roughly two-thirds of the time. When comparing GPT-4 and Claude, Claude is almost twice as likely to rank as second (behind the human-written story) and was the most preferred on three of the four assessments in which the New Yorker story was not chosen as the most favored. These ranking results confirm and accentuate the observation from the test passing rates analysis that Claude V1.3 generates higher-quality short stories than models in the GPT family.
@@
The attribution results paint a similar picture, with New Yorker stories predominantly attributed to an experienced writer, while LLM-generated stories get attributed to AI or an amateur writer. Interestingly, Claude V1.3 is more likely to be attributed to an amateur writer than an AI, whereas GPT3.5 and GPT4 stories are 80%+ attributed to AI. One hypothesis for such behavior could be that the participants in our study might be more familiar with text written by OpenAI models, as these models are commercially more successful, providing an element of surprise to Claude-generated text.
@@ 5.5. What can we infer from expert explanations of administering the TTCW?
Our annotation effort in Section 5.2.2 required experts to not only annotate for binary labels but also provide a justification paragraph accounting for their assessment. In Table 6, we provide examples of such justification for one of the TTCW tests for Originality. To gain insights into the main justifications experts provide for a story to pass or fail a TTCW test, we performed a manual thematic analysis of the expert justifications. We organized the results into a set of minimal phrases that often appear when a story passes or fails each TTCW test. Recent work has shown the utility of LLMs in clustering (Viswanathan et al., 2023). Based on these findings we asked the GPT4 model to cluster explanations across a given TTCW dimension into recurrent and broader representative themes. Three authors of the paper then manually verified these themes to ensure correctness. The outcome is summarized in Table 7 for Fluency and Flexibility tests, and Table 13 in the Appendix for Originality and Elaboration tests. The underlying themes found across the explanations reaffirm prior findings from Ippolito et al. (2022) where writers found LLM-generated stories experiencing difficulty in maintaining a style/voice and easily reverting to tropes and repetition as well as those from Mirowski et al. (2023) where screenwriters complained about the lack of subtext and character motivation.
E5
E3
E4
E1
E2
images, metaphors, etc was always very simple.
In an optional exchange with the expert participants (Section  5.3) who participated in the annotation (Section  5.2.2), they were given the opportunity to describe how they differentiated between AI-generated and human-written stories. Collected responses showed that expert evaluators did not make a decision on which stories were AI-generated based on factors such as grammatical characteristics. Table 8 lists the replies of a few experts, which we color-coded to highlight the recurrent issues that led them to believe that a story is AI-generated. Feedback was often aligned with individual TTCW, demonstrating that experts discriminated AI vs human written stories based on creative execution rather than spurious cues.
In particular, E5, E4, E1 thought AI struggles at Narrative Ending. E5 and E4 highlighted that AI-generated stories would forestall the ending by getting bigger in scope. E1 highlighted that AI-generated stories would have multiple disparate endings. E5, E4, E1, and E2 all highlighted that AI-generated stories would often contain abstruse and incoherent metaphors that do not add meaning or extremely cliched or simple metaphors thereby demonstrating poor Language Proficiency and use of Literary Devices. E1 highlighted one such example in a story - However, she managed to laugh louder and louder until her laughter transformed into an embrace of the sun’s atmosphere.
E1 and E2 further highlighted that characters in AI-generated stories have poor Rhetorical Complexity and are often lacking in subtext. E1 further added that AI-generated stories operate in a nearly opposite and Drax-like fashion in which there is only literal meaning, and that literal meaning is often nonsensical, or at least presented without any of the context that might make it seem like something a human would say or do. E1 highlighted an exchange below from an AI-generated story:
6. TTCW Implementation with LLMs as Assessors
Expert annotation such as the one we perform in Section 5 is costly: based on our evaluation protocol, evaluating a 1500-2500 word short story with a qualified expert costs \$20, and requires roughly 30 minutes of the expert’s time. Prior work has shown the promise of using LLMs in text evaluation. For instance, GPT3.5 and GPT4 are effective at evaluating the factual consistency of a summary to its document (Laban et al., 2023), or measuring the coherence of a summary (Gao et al., 2023). GPTEval (Liu et al., 2023) employs the framework of using large language models with chain-of-thoughts (CoT) (Wei et al., 2022) to assess the quality of NLG outputs. Recent work has also applied LLM-based evaluation to the creative domain (Rajani et al., 2023), claiming that GPT4 can achieve a high correlation with humans when evaluating brainstorming or creative generation tasks. In this section, we describe our TTCW implementation with LLMs as assessors to understand LLM’s ability to assess creative writing.
Prior work Wei et al. (2022) has shown how generating a chain of thought – a series of intermediate reasoning steps – significantly improves the ability of large language models to perform complex reasoning. Taking advantage of this we design the prompts/instructions for large language models in a slightly different fashion than for the human experts as can be seen in Table 3(Row 3 (Human Instruction) vs Row 4 LLM Instruction). To help the model make an informed decision we first ask it to list out elements specific to any given test such as “elements in the story that call to each of the five senses” for the World Building and setting test followed by asking it to decide overall and then provide its reasoning before choosing an answer between ‘Yes’ or ‘No’. The exact prompt contains (1) the story, (2) the expanded TTCW context, (3) the TTCW question, and (4) an LLM-specific instruction guiding the model to perform the task in a chain-of-though manner. We then measure model agreement with the majority vote of the three experts that conducted the test, using Cohen’s Kappa.
Dimension Test GPT3.5 GPT4 Claude
Fluency Understandability & Coherence -0.01 -0.01 -0.17
Narrative Pacing 0.05 0.0 -0.22
Scene vs Exposition -0.03 -0.08 -0.23
Literary Devices & Language Proficiency
0.04 -0.09 -0.11
Narrative Ending -0.02 0.02 0.02
Flexibility Emotional Flexibility -0.04 0.0 0.09
Perspective & Voice Flexibility 0.0 0.26 0.14
Structural Flexibility -0.04 0.0 -0.07
Originality Originality in Form 0.08 0.09 0.03
Originality in Thought 0.19 0.31 0.15
Originality in Theme & Content 0.06 -0.01 0.18
Elaboration World Building & Setting 0.0 0.00 0.09
Character Development -0.08 0.02 0.00
Rhetorical Complexity 0.0 0.0 0.02
Average 0.016 0.035 -0.006
6.1. Results
7. Discussion
7.1. Are experts good at detecting AI written texts or are they simply good at detecting lower quality texts?
During our evaluation, expert annotators were tasked with predicting a story’s author from three categories (Written by an Expert, Written by an Amateur, or AI-generated), even though none of the stories they annotated were selected from amateur authors. This mismatch was intended to reduce the likelihood of annotators tracing back the creation of the dataset and provide a more granular scale for their prediction.
Our goal is to design a rubric for creative writing evaluation that rates higher-quality text as better compared to lower-quality text. It is by no means designed to penalize AI-generated text. The primary reason for this was that it is really hard to detect AI-generated text because current large language models are good at generating fluent and human-like text. Certain traits are common in fiction written by an AI and an amateur writer which makes it difficult for the task of authorship attribution. One of our experts remarked There is, at best, little subtext in an amateur-written story, and likely none at all. This is the aspect that sometimes made me hesitate as to whether a story was written by AI or an amateur writer. While both AI and amateur-written fiction gravitate towards clichés, several experts mentioned AI writing tics that led them to believe it is not written by a human. Unlike text written by a mediocre writer, AI-written text would often contain predictable formations: strong topic sentences at the top of paragraphs; and summary sentences at the end of paragraphs; Some experts also mentioned peculiar sentence construction that was common in AI writing such as [time-orienting clause][comma][exposition]. Certain LLMs such as GPT4 also conflate good writing with ornamental use of language. This leads to generated text that is full of lofty superficial figurative language unlike text written by amateurs. Finally, AI’s voice in writing tended to be overly moralizing, unlike amateur writers. This evidence somewhat makes us believe that experts might have an easier time discriminating when lower-quality text is produced by a human vs a machine.
7.2. LLMs: from study subject to HCI research tool
This work leveraged LLMs as a tool to facilitate research, such as generating the first pass of the “expanded expert measure” text, or clustering expert-written explanations. Recent work has defined researchers and model developers as a main user category of LLMs (Chen et al., 2023a), with recent efforts for example making use of models to generate synthetic HCI research data (Hämäläinen et al., 2023). We reflect on the utility and limitations of using LLMs as a research tool in our work. We used an LLM in Section 5.2.1 to assist in writing the “expanded expert measure” descriptions, filling in the missing detail and context in expert-written descriptions, by adding definitions of technical terms or providing context or an example. LLMs were also useful in situations where we had to process a large volume of textual data, such as the clustering of the 2016 expert-written explanations we performed in Section 5.5. Manually clustering such explanations would have been prohibitively time-consuming, and LLMs enabled us to swiftly extract initial insights, which we discussed and refined. In either situation, we approach the use of LLMs in a human-AI collaborative framework, where the output of the LLM is validated by humans (in this case, the authors of the paper) and serves as an initial step in the research process rather than the final product.
7.3. Towards interactive LLM-based creative co-writing
In particular, we envision LLMs assisting Planning and Reviewing, crucial phases in the cognitive process theory of writing (Flower and Hayes, 1981). In prior work Gero et al. (2023) states that writers expressed the importance of specificity in the feedback instead of generic feedback like this might be a bit boring. We hope that the TTCW tests can provide the structure in future work looking to provide targeted feedback on writing. Further Ippolito et al. (2022) recently pointed out that professional writers constantly felt that LLM-generated text is rife with cliches and overused tropes. Metrics that quantify elements like originality in theme, structural flexibility, or rhetorical complexity could guide creative writing support tools [^11] https://www.sudowrite.com/ built using current LLMs, thereby improving planning and translation (Chung et al., 2021).
7.4. Is TTCW universal in nature?
The development of TTCW was informed by the expertise of 8 field specialists, with each test reflecting the insights of 1-3 experts. However, TTCW cannot be considered a universal benchmark for creative writing due to its reliance on a narrow expert base, potentially echoing Western literary biases due to the panel’s background in ”highbrow” literary fiction. This might marginalize other cultural narratives and styles. Specifically, TTCW evaluates creative writing through various lenses: TTCW Flexibility1 values diverse narrative perspectives, potentially disadvantaging stories focused on a singular viewpoint. TTCW Fluency2 and TTCW Flexibility2 assess the balance in storytelling elements, which might not favor experimental works that blend narrative techniques. Despite focusing on short stories, TTCW also critiques story closure in TTCW Fluency4, contrasting with plays that seek non-cathartic endings for specific impacts, as seen in the Brechtian tradition. Moreover, TTCW Originality3 rewards formal innovation, possibly underrating stories that creatively navigate within strict formal boundaries, like those adhering to traditional mythic structures. Thus, while TTCW offers a structured approach to evaluating creative writing, its scope and applicability are affected by certain choices and specific literary conventions it prioritizes.
8. Limitations and Future Work
Response Diversity and Temperature.
Coverage of the TTCW
Prompt Engineering and Optimization.
We open-source the prompts used to instruct the LLMs to generate the stories as well as prompts to administer the TTCW tests, which we iterated on and reviewed carefully. While there is no upper bound on engineering the ‘best’ prompt for a particular task, the output of an LLM is still dependent on the input prompt (Zhou et al., 2022). The LLMs that we evaluate in the study are closed-source models and the quality of their generation has been observed to change over time (Chen et al., 2023b). We hope future work can explore further refining of the prompts and parameters used in our work, and exploring their impact on the results.
Generalization beyond short fictional stories.
Objectivity in the Evaluation of Creativity.
9. Conclusion
Appendix A Appendix
A.1. NewYorker Data for evaluation
Figure 5. Distribution of word count of stories in our test set
Table 11 shows the data used for conducting our evaluation. The 12 stories shown are taken from The New Yorker and summarized into single-sentence plots. These stories come from highly established literary experts acting as an upper bound for what it means to be creative. These stories span complex themes.
story, do not say things like ‘Here’s the story […]:
Current story: {{story}}
Story Plot
A Triangle
a discarded, burned chair they’d noticed earlier.
Barbara Detroit,1966
Beyond Nature
Certain European Movies
Keys
way to reconnect with his present relationship through canoeing.
Listening For the Click
betrayal, love, and personal growth.
Maintenance, Hvidovre
Returns
a sunstroke, finding remnants of their last visit.
The Facade Renovation That’s Going Well
The Kingdom That Failed
Trash
The Last Dance with my Dad
A.2. Expert Perception on the TTCW tests
E5
structured way!
E3
E4
voice & diction!
E2
teasing out a fuller picture of things
E1
evaluate and appreciate literary writing.